# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v2

@@ 5. Numerical simulations
In this section, we illustrate the asymptotic stability of the origin w.r.t. the MPC closed loop using the kernel-EDMD surrogate in the prediction and optimization step by conducting numerical simulations. In the following, we denote by physics-informed kEDMD (PI-kEDMD) the models obtained with the modified Step 1 for the encoding of the equilibrium as described in Section 4.2, and with kEDMD the models obtained according to Bold et al. (2025b) without modifying Step 1.
5.1. Van der Pol
The first example is the Euler discretization of the nonlinear van-der-Pol oscillator. Due to using the Euler discretization, this directly yields the control-affine system
@@ 5.1. Van der Pol
with parameters $\Delta t=0.05$, $\nu=0.1$, and control constraints $\mathbb{U}=[-2,2]$ on the domain $\Omega=[-2,2]^{2}$. The kEDMD surrogate (21) is constructed using Wendland kernels with smoothness degree $k=1$. The cluster points are chosen according to a Padua grid. The set of Padua points of order $p$ in the domain $1,1]\times[-1,1]$ is defined as
@@
Moreover, we add a cluster point in the origin to the Padua grid, according to Assumption 3. The Padua grid is properly scaled to cover the domain $\Omega$. Padua grids of order $p\in\{25,50\}$ are considered, resulting in $d\in\{352,1327\}$ state-space sample points. In comparison to a uniform grid, the use of a Padua grid allows to have a more uniform error through the domain, and lower condition numbers in the model matrices. In fact, Padua points minimize, analogously to Chebychev nodes in one dimension, the Lebesgue constant and, thus, the condition number and are, thus, preferable. The use of a uniform grid with a large number of data points may lead to numerical problems in the solution of the optimal control problem.
@@
For each cluster point $x_{i}\in\mathcal{X}$, $d_{i}=25$ random control values $u_{ij}\in\mathbb{U}$ are chosen, yielding the data points $(x_{ij},u_{ij},x_{ij}^{+})$. Herein, $x_{ij}$ and $u_{ij}$ are drawn according to our data requirements with $r_{\mathcal{X}}=\nicefrac{{\sqrt{2}}}{{d}}$ to specify the neighborhood of $x_{i}$, see Section 4. For the derivation of the PI-kEDMD model Step 1 of the algorithm is modified for the cluster point $x_{1}=0$ by setting $\tilde{g}_{0}(x_{1})=0$, while in the kEDMD model Step 1 is performed in the same way for all cluster points. The MPC cost function is chosen as in (4) with matrices $Q=I_{2}$ and $R=10^{-4}$. The following closed-loop simulations emanate from the initial state $x^{0}=(0.5,0.5)^{\top}$, and consider a prediction horizon $N\in\{10,30\}$ in MPC. The simulations are implemented in Python, and Casadi Andersson et al. (2019) is used for the solution of the MPC optimization problem. The comparison of the closed loop error with different number of clusters and different prediction horizons is reported in Fig. 1.
@@
Figure 1. Van der Pol: Error $\|x(k)\|$ for the kEDMD-based MPC closed loop with different horizons and number of clusters.
@@
In the simulations with the kEDMD model, we observe practical asymptotic stability only: the error stagnates at a positive constant value. The tracking accuracy increases with the number of cluster points. This is expected, as a large number of cluster points is related to lower modeling errors. On the other hand, the PI-kEDMD models allow to reach a lower error, with a decrease up to the solver tolerance at around $10^{-14}$, in accordance to the shown asymptotic stability property of the MPC closed loop. We also notice that a fast decrease to the origin is obtained with the prolonged prediction horizon $N=30$ – a property related to the degree of suboptimality $\alpha_{N}$. For both the values of the prediction horizon, the asymptotic behavior of the PI-kEDMD-based MPC is the same that is obtained in a nominal MPC using the real system equations in the optimal control problem.
Figure 2. Van der Pol: trajectories of the kEDMD-based MPC closed loop with different horizons and number of clusters.
In Fig. 2, we report the trajectories of the same simulations in the phase space. We can see that, while the asymptotic behavior is strongly related to the accuracy of the model in the origin, the transient behavior is mainly influenced by the prediction horizon and the number of cluster points. Finally, Table 1 reports the average computation time for the solution of the MPC on a laptop with an Intel i7-1065G7 CPU. The computation times increase when considering larger numbers of cluster point and longer prediction horizons, and are comparable for the PI-kEDMD and the kEDMD model.
@@
Table 1. Average computation time in seconds of the kEDMD-based MPC for the Van der Pol oscillator, for different numbers of prediction horizons and of number of data points.
5.2. Four-tanks process
As a second example, we consider the four-tank system (see Alvarado et al. (2011)) described by the differential equation
@@ 5.2. Four-tanks process
The state of the system is given by the levels in the four tanks, i.e. $x=(h_{1},\;h_{2},\;h_{3},\;h_{4})^{\top}\in\mathbb{R}^{4}$, while the control variables are the flows through the two valves, i.e. $u=(q_{a},\;q_{b})^{\top}\in\mathbb{R}^{2}$. The numerical values of the system parameters can be found in Alvarado et al. (2011) and as sampling time we choose $\Delta t=10s$. The discrete-time version of (28) obtained with the forward Euler method is considered as ground truth. The controlled state for this system is the equilibrium given by $\bar{x}=(0.65,0.66,0.6417,0.6882)^{\top}m$ and $\bar{u}=(1.666,1.974)^{\top}m^{3}/h$. To derive the kEDMD surrogate model, the state and inputs are shifted around the origin. The state domain, where we sample, is $\Omega=[0.2\ m,1.36\ m]^{2}\times[0.2\ m,1.30\ m]^{2}$, while the input is constrained by the set $\mathbb{U}=[0\ m^{3}/h,3.26\ m^{3}/h]\times[0\ m^{3}/h,4\ m^{3}/h]$. The cluster centers are chosen in a uniform grid in $\Omega$. Moreover, an additional cluster center is added at the equilibrium. We consider datasets with $d\in\{626,1297,2402\}$ cluster centers, that correspond to $\sqrt[4]{d-1}\in\{5,6,7\}$. We decided to stick to a uniform grid, because the explicit construction of a grid minimizing the Lebesgue constant in dimensions higher than two is more involved and we did not run into numerical issues. Possible approaches to reduce the condition number of the model matrices are the use of the cartesian product of Padua grids or of more sophisticated approaches as outlined, e.g., in Jiménez-Ramos et al. (2025). The cluster radius is $r_{\mathcal{X}}=2/d$ for all the models, and 25 data points $(x_{ij},u_{ij},x_{ij}^{+})$ are considered in each cluster. For the closed-loop simulations, we consider an initial state $x_{0}=(1.0,1.0,1.0,1.0)^{\top}m$, MPC cost matrices $Q=I_{4}$ and $R=10^{-4}I_{2}$, and a prediction horizon $N=10$.
@@
PI-kEDMD KEDMD
$d=626$ 0.370 0.383
$d=1297$ 0.875 0.864
$d=2402$ 1.876 1.744
Table 2. Average computation time in seconds of the kEDMD-based MPC for the four-tank system, for different number of data points and prediction horizon $N=10$.

### v3

@@ 5. Numerical simulations
In this section, we illustrate the asymptotic stability of the origin w.r.t. the MPC closed loop using the kernel-EDMD surrogate in the prediction and optimization step by conducting numerical simulations. In the following, we denote by physics-informed kEDMD (PI-kEDMD) the models obtained with the modified Step 1 for the encoding of the equilibrium, and with kEDMD the models obtained according to Bold et al. (2025b) without modifying Step 1.
Van der Pol
We consider the Euler discretization of the nonlinear van-der-Pol oscillator. Due to using the Euler discretization, this directly yields the control-affine system
@@
with parameters $\Delta t=0.05$, $\nu=0.1$, control constraints $\mathbb{U}=[-2,2]$ and domain $\Omega=[-2,2]^{2}$. The kEDMD surrogate (21) is constructed using Wendland kernels with smoothness degree $k=1$. The cluster points are chosen according to a Padua grid. The set of Padua points of order $p$ in the domain $1,1]\times[-1,1]$ is defined as
@@
Moreover, we add a cluster point in the origin to the Padua grid, according to Assumption 3. The Padua grid is properly scaled to cover the domain $\Omega$. Padua grids of order $p\in\{25,50\}$ are considered, resulting in $d\in\{352,1327\}$ state-space sample points. In comparison to a uniform grid, the use of a Padua grid allows to have a more uniform error through the domain, and lower condition numbers in the model matrices. In fact, Padua points minimize, analogously to Chebychev nodes in one dimension, the Lebesgue constant and, thus, the condition number and are, thus, preferable. The use of a uniform grid with a large number of data points may lead to numerical problems in the solution of the optimal control problem. For each cluster point $x_{i}\in\mathcal{X}$, $d_{i}=25$ random control values $u_{ij}\in\mathbb{U}$ are chosen, yielding the data points $(x_{ij},u_{ij},x_{ij}^{+})$. Herein, $x_{ij}$ and $u_{ij}$ are drawn according to our data requirements with $r_{\mathcal{X}}=\nicefrac{{\sqrt{2}}}{{d}}$ to specify the neighborhood of $x_{i}$, see Section 4. For the derivation of the PI-kEDMD model Step 1 of the algorithm is modified for the cluster point $x_{1}=0$ by setting $\tilde{g}_{0}(x_{1})=0$, while in the kEDMD model Step 1 is performed in the same way for all cluster points. The MPC cost function is chosen as in (4) with matrices $Q=I_{2}$ and $R=10^{-4}$. The following closed-loop simulations emanate from the initial state $x^{0}=(0.5,0.5)^{\top}$, and consider a prediction horizon $N\in\{10,30\}$ in MPC. The simulations are implemented in Python, and Casadi is used for the solution of the MPC optimization problem. The comparison of the closed loop error with different number of clusters and different prediction horizons is reported in Fig. 1.
@@
Figure 1. Van der Pol: Error $\|x(k)\|$ for the kEDMD-based MPC closed loop (left) and trajectories with different horizons and number of clusters (right). Solid and dash-dotted lines are used for PI-kEDMD and dashed and dotted lines for kEDMD.
@@
In the simulations with the kEDMD model, we observe practical asymptotic stability only: the error stagnates at a positive constant value. The tracking accuracy increases with the number of cluster points. This is expected, as a large number of cluster points is related to lower modeling errors. Instead, the PI-kEDMD models allow to reach a lower error, with a decrease up to the solver tolerance at around $10^{-14}$, in accordance to the shown asymptotic stability property of the MPC closed loop. We also notice that a fast decrease to the origin is obtained with the prolonged prediction horizon $N=30$ – a property related to the degree of suboptimality $\alpha_{N}$. For both the values of the prediction horizon, the asymptotic behavior of the PI-kEDMD-based MPC is the same that is obtained in a nominal MPC using the real system equations in the optimal control problem. In Fig. 1 (right), we report the trajectories of the same simulations in the phase space. We can see that, while the asymptotic behavior is strongly related to the accuracy of the model in the origin, the transient behavior is mainly influenced by the prediction horizon and the number of cluster points. Finally, Table 1 reports the average computation time for the solution of the MPC on a laptop with an Intel i7-1065G7 CPU. The computation times increase when considering larger numbers of cluster point and longer prediction horizons, and are comparable for the PI-kEDMD and the kEDMD model.
@@
$N=10$ $N=10$
four-tanks $d=626$ 0.370 0.383
$d=1297$ 0.875 0.864
$d=2402$ 1.876 1.744
Table 1. Average computation time in seconds of the kEDMD-based MPC for the Van der Pol oscillator (V.d.P.) and for the four-tanks process, for different prediction horizons and numbers of data points.
Four-tanks process
Our second example is the four-tank system Alvarado et al. (2011) described by the differential equation
@@
The state of the system is given by the levels in the four tanks, i.e. $x=(h_{1},\;h_{2},\;h_{3},\;h_{4})^{\top}\in\mathbb{R}^{4}$, while the control variables are the flows through the two valves, i.e. $u=(q_{a},\;q_{b})^{\top}\in\mathbb{R}^{2}$. The numerical values of the system parameters can be found in Alvarado et al. (2011) and as sampling time we choose $\Delta t=10s$. The discrete-time version of (28) obtained with the forward Euler method is considered as ground truth. The controlled state for this system is the equilibrium given by $\bar{x}=(0.65,0.66,0.6417,0.6882)^{\top}m$ and $\bar{u}=(1.666,1.974)^{\top}m^{3}/h$. To derive the kEDMD surrogate model, the state and inputs are shifted around the origin. The state domain, where we sample, is $\Omega=[0.2\ m,1.36\ m]^{2}\times[0.2\ m,1.30\ m]^{2}$, while the input is constrained by the set $\mathbb{U}=[0\ m^{3}/h,3.26\ m^{3}/h]\times[0\ m^{3}/h,4\ m^{3}/h]$. The cluster centers are chosen in a uniform grid in $\Omega$. Moreover, an additional cluster center is added at the equilibrium. We consider datasets with $d\in\{626,1297,2402\}$ cluster centers, that correspond to $\sqrt[4]{d-1}\in\{5,6,7\}$. We decided to stick to a uniform grid, because the explicit construction of a grid minimizing the Lebesgue constant in dimensions higher than two is more involved and we did not run into numerical issues. Possible approaches to reduce the condition number of the model matrices are the use of the cartesian product of Padua grids or of more sophisticated approaches as outlined, e.g., in Jiménez-Ramos et al. (2025). The cluster radius is $r_{\mathcal{X}}=2/d$ for all the models with 25 samples $(x_{ij},u_{ij},x_{ij}^{+})$ in each cluster. For the closed-loop simulations, we consider an initial state $x_{0}=(1.0,1.0,1.0,1.0)^{\top}m$, MPC cost matrices $Q=I_{4}$ and $R=10^{-4}I_{2}$, and a prediction horizon $N=10$.